Session: Feb 1998 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Wed Feb 18 10:46:25 MET 1998) Minor failures
The schedule did not allow time for tape change, thus two scanes (beginning at 22:36) were lost. Also lost is scan at 23:03 because the source fell outside the telescope limits.

Noto (C.Trigilio at Fri Feb 20 15:48:59 MET 1998) Minor failures
Serious power failure at 10:55 : VLBA terminal and computers down. Re-boot of everything. Observations restart at 11:35. Scan 12:25 lost. Tape change: snap file and list file indicate only one tape, but two tape labels printed; observation consistent with TWO tapes. NEWTAPE inserted manually at 22:45 - DRUDG BUG?!?

Calibration (jean\-francois Desmurs at Mon Mar 2 18:47:27 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name gl022.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Thu Mar 5 15:56:51 MET 1998) Minor failures
0922-1000 source below elv limit. 21 seconds for tape change, so lost some data on change. NGC7682 below elv limit.

Effelsberg (Dharam Vir Lal at Fri Mar 6 06:34:00 MET 1998) Minor failures